OpenStudy (anonymous):

What is the half-life of a radioactive substance if it takes 5 years for one-third of the material to decay?

OpenStudy (anonymous):

If I use the formula P(t) = Ce^rt, I will just plug in to get 1/3 = e^5r and then solve for r. I got r = ln |1/3| /5 and then i plug back into the equation to get half life and solve for time?? 1/2 = e^(ln |1/3|/5)t and solving for t i get 5 ln |1/2|/ ln |1/3| = t and then i am done???

OpenStudy (anonymous):

is this correct?

OpenStudy (kropot72):

\[\frac{1}{3}=e ^{-k5}\] Solve to find the decay constant k. The plug in the value of k into the following and solve for t.\[\frac{1}{2}=e ^{-kt}\]

OpenStudy (anonymous):

y did you use -k? is that part of the formula, i thought it was just a regular constant.

OpenStudy (anonymous):

so k = -5/ln |1/3|

OpenStudy (anonymous):

and t = (ln |1/2| ln |1/3| )/ 5 can you simplify the top

OpenStudy (kropot72):

\[k=\frac{\ln \frac{1}{3}}{-5}\]

OpenStudy (kropot72):

k=0.219722457

OpenStudy (anonymous):

dang it. okay i see my mess up on k

OpenStudy (kropot72):

In the formula for radioactive decay the exponent of e is negative.

OpenStudy (anonymous):

so t will be 5(ln |1/2|)/ln |1/3|

OpenStudy (anonymous):

can we simplify that?

OpenStudy (kropot72):

\[Half\ life=\frac{\ln \frac{1}{2}}{-0.219722457}\]

OpenStudy (kropot72):

No need to simplify. Just use a calculator.

OpenStudy (anonymous):

so t = 3.15 years?

OpenStudy (kropot72):

Correct.

OpenStudy (anonymous):

but if it takes 5 years for 1/3, then how does that make sense???

OpenStudy (kropot72):

Actually I misread the question. The amount of material left after 5 years is 2/3 of the original amount. So recalculating: \[\ln \frac{2}{3}=-5k\] \[k=\frac{\ln \frac{2}{3}}{-5}=0.081\] \[half\ life=\frac{\ln \frac{1}{2}}{-0.081}\]

OpenStudy (anonymous):

Okay, so you use the amout left and not the amount used. makes much more sense now... thank you.

OpenStudy (kropot72):

You're welcome :)